OpenStudy (yolo_king):

A card is chosen at random from a deck of 52 cards. It is replaced and a second card is chosen. What is the probability both cards will be an Ace? A. 1/169 B. 2/13 C. 1/26 D. 1/13

OpenStudy (yolo_king):

Can someone teach me the concept?

OpenStudy (yolo_king):

I dont understand how many aces are there in the deck of 52?????

OpenStudy (sweetburger):

Ok so you have 4 aces out of a deck of 52 cards. The probability of getting an ace card is 4 aces/52 total cards.

OpenStudy (yolo_king):

Oh ok so 1/13

OpenStudy (sweetburger):

Now thats not what the question is asking. The question wants to know what would happen if you on your first draw pick an ace card then put back the ace card and on your second draw pick another ace card. This can be represented by \[\frac{ 4 }{ 52 }\times \frac{ (3+1) }{ (51+1) }\] written like this it stresses that the missing ace card was placed back into the deck. However, it is not necessary to write it like this you can just instead say \[\frac{ 4 }{ 52 }\times \frac{ 4 }{ 52 }\]

OpenStudy (yolo_king):

OOpps i mean 1/169 cause 4/52 * 4/52 = 16/2704 simplified would be 1/169

OpenStudy (sweetburger):

Yes, correct!

OpenStudy (yolo_king):

Omgird thanks a lot!!! Couple more?

OpenStudy (sweetburger):

sure ill see if i can help. just open a new question